# PSLM: Parallel Generation of Text and Speech with LLMs for Low-Latency Spoken Dialogue Systems

### v1

@@ 2.2 Integrating LMs with a Speech Stream
PSLM is built on top of a pretrained decoder-only Transformer (Vaswani et al., 2017). We add new input embedding and output projection layers to process speech tokens, while the structure of the intermediate Transformer layers remains unchanged. We randomly initialize the weights of new embedding and projection layers.
A challenge of joint text-speech modeling lies in the mismatch in their lengths. In this study, we simply right-pad TQ and TA sequences with special [TEXT-PAD] tokens to align their lengths with those of the SQ and SA sequences, respectively. With this approach, we force the PSLM to learn alignment between text and speech. Our PSLM is trained by minimizing the sum of cross entropy losses for each stream. We include prompt tokens, comprising TQ and SQ, in the loss calculation. During inference, PSLM receives these prompt tokens and generates TA and SA in parallel.
@@ 2.3 Introducing Multiple Speech Streams
For further acceleration, we introduce multiple speech streams to PSLM. Assume that PSLM has $1+S$ streams, one for text tokens and $S$ for speech tokens. We first reshape the original speech token sequence of length $N$ into a matrix with dimensions $(N/S,S)$, and use the $s$-th column as the $s$-th speech stream ($s\in\{1,\ldots,S\})$. This reduces the sequence length to $1/S$, which leads to approximately an $S$ times speedup. During training, simply summing the cross entropy losses for each stream makes the loss of text tokens less dominant, leading to poor text generation quality. Therefore, we introduce a weighted loss, where we multiply the loss for speech streams by $1/S$ to balance the weight of losses for text and speech streams.
@@ 2.4 Streaming Inference with HiFi-GAN
Following Chen et al. (2022), we use HiFi-GAN for streaming inference; specifically, we provide partial speech tokens to generate waveform fragments. In this study, we use non-causal convolution to maintain high speech quality. Therefore, the first speech fragment can be generated once $N_{\textrm{offset}}=\lfloor R/2\rfloor+1$ tokens are decoded, where $R$ denotes the receptive field of HiFi-GAN. Implementation details can be found in Appendix 2.
@@ 3.1 Dataset
We used an internal dataset comprising 1.8M written QA pairs for training and a public Japanese instruction dataset (Hayashibe, 2023) comprising 669 written QA pairs for evaluation. We filtered out samples with more than 140 characters in TA, obtaining 396 samples for evaluation. For both sets, we constructed a spoken question answering (SQA) dataset by synthesizing SQ and SA using a well-trained single-speaker TTS system based on VITS (Kim et al., 2021).
@@ 3.3 Baselines
We involved three CoM-based baselines, which share the model weights but differ in their prompts during decoding: (1) CoM-SQ receives only SQ, (2) CoM-ASR receives SQ and transcribed TQ, and (3) CoM receives SQ and gold TQ.
@@ 3.4 Evaluation Metrics
We simulated latency according to Equations 2 and 3 for each sample in the evaluation set, and reported the median values. Further details can be found in Appendix E.1.
@@
We also conducted two subjective evaluations: one for text and the other for speech. In the text evaluation, we presented pairs of gold TQ and generated TA, along with the criteria used in ChatGPT-based evaluation. In the speech evaluation, we presented gold SQ and generated SA successively, along with their TQ and TA, and asked the raters to evaluate (1) how natural the SA is as the speech of the TA, and (2) whether the response is fast enough. The duration of silence between SQ and SA was simulated in the manner described in Section 2.5, except for the Ground Truth where the silence duration was set to 200ms, the average turn-taking gap in human conversation (Levinson and Torreira, 2015). Scores were rated on a 5-point scale. Fifty samples were randomly chosen from the evaluation set, and twenty in-house workers rated twenty samples each.
@@ 4.1 Automatic Evaluation
To answer RQ2, we trained PSLM variants with two (-2x) or three (-3x) speech streams [^5] PSLM-3x was trained with a batch size of 4 due to the increased number of parameters.. PSLM-2x achieved comparable scores to PSLM, whereas PSLM-3x demonstrated significant degradation. From these results, we conclude that speech tokens can be decoded in up to two streams without quality degradation (RQ2). An ablation study can be found in Appendix D.
@@ 4.2 Human Evaluation
Considering practical applicability to SQA, we manually evaluated three methods: CoM-SQ, CoM-ASR, and PSLM-ASR, which do not rely on gold TQ, along with Ground-Truth. Table 2 shows the results. The text response naturalness of PSLM-ASR was comparable to CoM-ASR and higher than CoM-SQ, which is consistent with the automatic evaluation results. For speech naturalness, all methods achieved higher scores than Ground-Truth. This result can be attributed to two reasons: (1) SA of Ground-Truth are synthetic speech, which may include errors in pronunciation, intonation, and pauses, and (2) SA of Ground-Truth are typically longer than those of other methods, incurring that one or two unnatural parts lowered the entire score. Nevertheless, we confirmed that our approach can generate natural and faithful speech responses. For response speed evaluation, PSLM-ASR achieved a significantly higher score than CoM-ASR and CoM-SQ. This finding verifies that the proposed method reduces latency both numerically and perceptibly. Detailed analysis can be found in Appendix E.2.
@@ 6 Limitations
We recognize several limitations of this study. First, PSLM sacrifices ASR capability for faster response, requiring an external ASR module to serve as a spoken dialogue system. Enabling ASR with the PSLM architecture can be an interesting research direction. Second, we used single-speaker synthetic speech for SQ and SA, which lacks diversity in several aspects of speech such as accent, rhythm, emotion, and timbre. Practical applications may require to accept voices of arbitrary speakers, which we will address in future work. Finally, multi-turn dialogue settings were not investigated in our experiments. While SpeechGPT (Zhang et al., 2023) was not applied to multi-turn dialogue due to sequence length limitations, our models with multiple speech streams have the potential to perform multi-turn dialogue.
@@ Appendix C ChatGPT Evaluation Prompt
Appendix D Ablation Study
@@
Appendix E Latency Analysis
E.1 Configuration for Latency Simulation
For latency calculation, we set $D_{\textrm{s2t}}=0.05$, $D_{\textrm{SQ}}=0.05$, $D_{\textrm{ASR}}=0.2$, and $D_{\textrm{t2s}}=0.01$ based on measurements on a single NVIDIA A100 GPU. The actual TPS varies with resources and optimization; 70 TPS was achieved with vLLM (Kwon et al., 2023) optimization, and 25 TPS without it. Meanwhile, to enable streaming inference with HiFi-GAN, LMs need to generate 50 speech tokens per second. Thus, we set $P$ to 50 for Table 1 and 50 or 100 in the next section.
E.2 Detailed Latency Analysis
We investigated the overall latency for different TA lengths. Figure 4 shows the results. Due to the need for prior generation of TA, the latency of CoM-SQ and CoM-ASR increases linearly as TA length increases. In contrast, the latency of PSLM-ASR is constant because Equation 3 does not include $N_{\textrm{TA}}$, and PSLM-2x-ASR further reduces the latency. The gap between CoM-based and PSLM-based systems is remarkable when generating long TA, highlighting the effectiveness of generating text and speech tokens in parallel.
Figure 3: A prompt for ChatGPT evaluation.
@@ E.2 Detailed Latency Analysis
Figure 4: Latency vs. TA length for different methods and token generation speed. PSLM-2x-ASR (50 TPS) is omitted because its latency is identical to PSLM-ASR (100 TPS).

### v2

@@ 2.2 Integrating LMs with a Speech Stream
PSLM is built on top of a pretrained decoder-only Transformer (Vaswani et al., 2017). An overview of the PSLM architecture is provided in Figure 2. We add new input embedding and output projection layers to process speech tokens, while the structure of the intermediate Transformer layers remains unchanged. The embeddings of text and speech tokens are summed before being fed to the Transformer layers. The hidden features from the final Transformer layer are passed to two output projection layers to calculate the logits of the next text and speech tokens. We randomly initialize the weights of new embedding and projection layers.
A challenge of joint text-speech modeling lies in the mismatch in their lengths. In this study, we simply right-pad TQ and TA sequences with special [TEXT-PAD] tokens to align their lengths with those of the SQ and SA sequences, respectively. In a preliminary experiment on the CoM-based architecture, we attempted to generate text tokens and their corresponding speech tokens alternatively in a similar manner to ELLA-V (Song et al., 2024); however, this approach led to frequent mispronunciation. This is mainly because, in our case, the text is represented by tokens rather than phonemes; in some languages, the pronunciation of a character often changes according to subsequent characters, and a certain amount of lookahead is necessary to achieve accurate pronunciation. In contrast, our alignment strategy allows the model to focus on text token generation initially and then refer to the generated text when producing the majority of speech tokens, leading to more accurate pronunciation.
Our PSLM is trained by minimizing the sum of cross entropy losses for each stream. We include prompt tokens, comprising TQ and SQ, in the loss calculation. During inference, PSLM receives these prompt tokens and generates TA and SA in parallel. Text and speech tokens are sampled independently from their respective distributions.
Figure 2: Architecture of PSLM.
@@ 2.3 Introducing Multiple Speech Streams
For further acceleration, we introduce multiple speech streams to PSLM. Assume that PSLM has $1+S$ streams, one for text tokens and $S$ for speech tokens. Given the original speech token sequence of length $N$, the $s$-th speech stream consists of the speech tokens with indices $s,s+S,s+2S,...,s+MS$, where $s\in\{1,\ldots,S\}$ and $M=\lfloor N/S\rfloor-1$. Compared to simply increasing the batch size, where the system’s throughput improves but the latency for each instance remains unchanged, our approach reduces the sequence length handled by the Transformer layers to $1/S$, leading to an approximate $S$-fold speedup even in the single-instance scenario.
During training, simply summing the cross entropy losses for each stream makes the loss of text tokens less dominant, leading to poor text generation quality. Therefore, we introduce a weighted loss, where we multiply the loss for speech streams by $1/S$ to balance the weight of losses for text and speech streams.
@@ 2.4 Streaming Inference with HiFi-GAN
Following Chen et al. (2022), we use HiFi-GAN for streaming inference; specifically, we provide partial speech tokens to generate waveform fragments. In this study, we use non-causal convolution to maintain high speech quality. Therefore, the first speech fragment can be generated once $N_{\textrm{offset}}=\lfloor R/2\rfloor+1$ tokens are decoded, where $R$ denotes the receptive field of HiFi-GAN. Implementation details can be found in Appendix B.
@@ 3.1 Dataset
We used an internal dataset comprising 1.8M written QA pairs for training all models. Since some of these samples, which were primarily crawled from the internet, were deemed unsuitable for evaluation, we used a publicly available Japanese dataset (Hayashibe, 2023) for evaluation. This dataset was manually reviewed and consists of 669 diverse written QA pairs. We further filtered the evaluation set by excluding samples whose TQ or TA exceeded 140 characters, the maximum number of characters observed in the training set. The final evaluation set contained 396 samples. For both the training and evaluation sets, we constructed a spoken question answering (SQA) dataset by synthesizing SQ and SA using a well-trained single-speaker TTS system based on VITS (Kim et al., 2021).
@@ 3.3 Baselines
We involved three CoM-based baselines, which share the model weights but differ in their prompts during decoding: (1) CoM-SQ receives only SQ, (2) CoM-ASR receives SQ and transcribed TQ, and (3) CoM receives SQ and gold TQ. In our preliminary experiments, the three-stage training (Zhang et al., 2023) was not effective in our configuration; thus, we trained the model using the same configuration as described in Section 3.2.
@@ 3.4 Evaluation Metrics
We simulated latency according to Equations 2 and 3 for each sample in the evaluation set, and reported the median values. We set $D_{\textrm{s2t}}=0.05$, $D_{\textrm{SQ}}=0.05$, $D_{\textrm{ASR}}=0.2$, and $D_{\textrm{t2s}}=0.01$ based on measurements taken on a single NVIDIA A100 GPU. For the TPS value $P$, the actual TPS varies depending on computing resources and optimization; 70 TPS was achieved with vLLM (Kwon et al., 2023) optimization, and 25 TPS without it. Meanwhile, for streaming inference with HiFi-GAN, LMs need to generate 50 speech tokens per second. Therefore, we set $P$ to 50 in our simulations to match this requirement.
@@
We also conducted two subjective evaluations: one for text and the other for speech. In the text evaluation, we presented pairs of gold TQ and generated TA, and raters evaluated the naturalness of TA based on the same criteria used in the ChatGPT-based evaluation (Text Naturalness). In the speech evaluation, we presented gold SQ and generated SA successively, along with their TQ and TA, and asked the raters to evaluate (1) how natural the SA is as the speech of the TA (Speech Naturalness), and (2) whether the response is fast enough (Speed Score). For better reproducibility, we provide the actual instruction used for speech evaluation in Appendix D. The duration of silence between SQ and SA was simulated in the manner described in Section 2.5, except for the Ground Truth where the silence duration was set to 200ms, the average turn-taking gap in human conversation (Levinson and Torreira, 2015). Scores were rated on a 5-point scale. Fifty samples were randomly chosen from the evaluation set, and twenty in-house workers rated twenty samples each.
@@ 4.1 Automatic Evaluation
To answer RQ2, we trained PSLM variants with two (-2x) or three (-3x) speech streams [^5] PSLM-3x was trained with a batch size of 4 due to the increased number of parameters.. PSLM-2x achieved comparable scores to PSLM, whereas PSLM-3x demonstrated significant degradation. From these results, we conclude that speech tokens can be decoded in up to two streams without quality degradation (RQ2). An ablation study can be found in Appendix E.
@@ 4.2 Human Evaluation
Considering practical applicability to SQA, we manually evaluated three methods: CoM-SQ, CoM-ASR, and PSLM-ASR, which do not rely on gold TQ, along with Ground-Truth. Table 2 shows the results. The text response naturalness of PSLM-ASR was comparable to CoM-ASR and higher than CoM-SQ, which is consistent with the automatic evaluation results. For speech naturalness, all methods achieved higher scores than Ground-Truth. This result can be attributed to two reasons: (1) SA of Ground-Truth are synthetic speech, which may include errors in pronunciation, intonation, and pauses, and (2) SA of Ground-Truth are typically longer than those of other methods, incurring that one or two unnatural parts lowered the entire score. Nevertheless, we confirmed that our approach can generate natural and faithful speech responses. For response speed evaluation, PSLM-ASR achieved a significantly higher score than CoM-ASR and CoM-SQ. This finding verifies that the proposed method reduces latency both numerically and perceptibly. Detailed analysis can be found in the next subsection.
@@
4.3 Detailed Latency Analysis
The sequence length of TA, or $N_{\textrm{TA}}$, is the most influential factor in overall latency of CoM-based systems, as TA must be generated before SA. Thus, we investigated the overall latency by varying $N_{\textrm{TA}}$. Figure 3 shows the results. Due to the need for prior generation of TA, the latency of CoM-SQ and CoM-ASR increases linearly as TA length increases. In contrast, the latency of PSLM-ASR is constant because Equation 3 does not include $N_{\textrm{TA}}$, and PSLM-2x-ASR further reduces the latency. The gap between CoM-based and PSLM-based systems is remarkable when generating long TA, highlighting the effectiveness of generating text and speech tokens in parallel.
Figure 3: Latency vs. TA length for different methods and tokens per second (TPS). PSLM-2x-ASR (50 TPS) is omitted because its latency is identical to PSLM-ASR (100 TPS).
@@ 6 Limitations
We recognize several limitations of this study. First, PSLM sacrifices ASR capability for faster response, requiring an external ASR module to serve as a spoken dialogue system. Although this dependency can complicate the system structure, it does not degrade the system’s performance, provided that an appropriate ASR module is selected. This is supported by the fact that CoM-ASR outperformed CoM-SQ, as described in Section 4.1. Nevertheless, enabling ASR with the PSLM architecture can be an interesting research direction. Second, we used single-speaker synthetic speech for SQ and SA, which lacks diversity in several aspects of speech such as accent, rhythm, emotion, and timbre. Practical applications may require to accept voices of arbitrary speakers, which we will address in future work. Finally, multi-turn dialogue settings were not investigated in our experiments. While SpeechGPT (Zhang et al., 2023) was not applied to multi-turn dialogue due to sequence length limitations, our models with multiple speech streams have the potential to perform multi-turn dialogue.
@@ Appendix C ChatGPT Evaluation Prompt
Appendix D Speech Evaluation Instruction
We used the instruction in Figure 6 for speech evaluation. The original instruction was written in Japanese, but a translated version is presented here.
Appendix E Ablation Study
@@
Figure 5: Prompt for ChatGPT evaluation.
Figure 6: Instruction for speech evaluation.